OpenStudy (anonymous):

Find the amount if $900 is invested at 15% compounded monthly for 15 months.

OpenStudy (jdoe0001):

hmm how many "years" are in 15months?

OpenStudy (anonymous):

1 year 3 months

OpenStudy (jdoe0001):

well.. yes... thus \(\bf 1\frac{1}{4}\to 1.25\ years\)

OpenStudy (jdoe0001):

\(\bf A=P\left(1+\frac{r}{n}\right)^{nt} \\ \quad \\ A=\textit{current amount}\\ P=\textit{original amount deposited}\to \$900\\ r=rate\to 15\%\to \frac{15}{100}\to 0.15\\ n=\textit{times it compounds per year, monthly thus}\to 12\\ t=years\to 1.25 \\ \quad \\ A=900\left(1+\frac{0.15}{15}\right)^{15\cdot 1.25}\)

OpenStudy (dumbcow):

@jdoe0001 , n=12 \[A = 900(1+ \frac{.15}{12})^{12*1.25}\]

OpenStudy (anonymous):

A=1084.346

OpenStudy (anonymous):

Thank you to you both for helping me figure out the formula.

OpenStudy (jdoe0001):

ohh shoot.. got a typo yeahh yes hehe \(\bf A=900\left(1+\frac{0.15}{12}\right)^{15\cdot 1.25}\)

OpenStudy (jdoe0001):

actually \(\bf A=900\left(1+\frac{0.15}{12}\right)^{12\cdot 1.25}\)

OpenStudy (anonymous):

Thank you!

OpenStudy (jdoe0001):

yw